A rail gun accelerates a projectile with a constant acceleration of 50.0 m/s2. The projectile starts at rest and moves1.00 m along the rail. Atthe end of the rail, what is the projectile’s speed?
1
Expert's answer
2020-04-21T11:01:43-0400

The final speed of a projectile

"v_f=\sqrt{v_i^2+2ad}\\
=\sqrt{0^2+2\times 50.0\times 1.00}=10\:\rm m/s."